Wildcat Ltd, a manufacturing company sold a machinery for Rs 8 lacs at the year end. The company had purchased the machinery four years back for Rs 15 lacs and had depreciated the same using written down value method of depreciation @ 20%. As an accounts executive of Wildcat Ltd, calculate the WDV of the asset for the four years, accumulated depreciation for four years and profit/loss on sale, if any.
1
Expert's answer
2019-11-11T15:35:25-0500

The WDV of the asset for the four years is:

Rate of depreciation = 1 - (8/15)^(1/4) = 0.145.

Accumulated depreciation for four years was: 15×0.2×4 = 12.

Profit on sale is 12 - 8 = 4.